OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of 1/3 and passes through the point (-6, 0). A.Y-0=-1/3(X+6) B.Y-0=1/3(X-6) C.Y+6=1/3(X-0) D.Y-6=1/3(X-0)

OpenStudy (anonymous):

Point-slope form: \[y - y_{1} = m(x - x_{1})\] Plug in 1/3 for m and the point for x_1 and y_1

OpenStudy (anonymous):

IS IT B

OpenStudy (anonymous):

You sure you wrote B correctly? It should be: \[y - 0 = \frac13(x + 6)\]

OpenStudy (anonymous):

Or did you forget a negative somewhere?

OpenStudy (anonymous):

yeah srry i meant d

OpenStudy (radar):

Using the equation provided by Calcmathlete: y-0=1/3(x-(-6) y-0=1/3(x+6) It aint B

OpenStudy (anonymous):

its d

OpenStudy (radar):

Think

OpenStudy (anonymous):

oh its a

OpenStudy (radar):

Yeah, I think you probably mishandled the sign.

OpenStudy (anonymous):

thank you

OpenStudy (radar):

Think some more......the slope must be a positive 1/3 To be honest, I don;t see the answer yet.

OpenStudy (radar):

Now I do, and the D was right, let me rearrange D Y=1/3x+6 Sorry you were correct, and I was the one that needed to think Lol

OpenStudy (radar):

Don't know why they put D. in such a way. But it turns out to be the best answer.

OpenStudy (anonymous):

so its D.

OpenStudy (radar):

Yes it is D. D when arranged as a "slope intercept" form is correct. They have tried to make it tricky.!

OpenStudy (anonymous):

OKAY THANK YOU

OpenStudy (radar):

Sorry, if I steered you away from D. It was my fault, I just did not imagine what it would look like if arranged properly.

OpenStudy (anonymous):

ITS FINE DONT WRRY ABOUT IT